OpenStudy (anonymous):

<θ is in standard position with its terminal arm in the stated quadrant, and 0≤θ<2pi. A trigonometric ratio: cosθ= -2/3 is given in quadrant III. How do I find the exact values of the other two trignometric ratios?

myininaya (myininaya):

There are 6 trig functions Do you mean sin(theta) and tan(theta)? Because there is also csc(theta), sec(theta), cot(theta)

OpenStudy (anonymous):

sin(theta) and tan(theta)

myininaya (myininaya):

I'm having trouble reading your inequality

myininaya (myininaya):

\[0 \le \theta \le 2 \pi ?\]

OpenStudy (anonymous):

sorry , i mean 0≤θ≤2pi.

OpenStudy (anonymous):

in quadrant III

myininaya (myininaya):

Well we know the triangle can either be in quadrant 2 or 3 since we want x to be negative since we have cos(theta)=-2/3 =(adj/hyp) hyp is positive

myininaya (myininaya):

oh its in quadrant 3 ok!

myininaya (myininaya):



myininaya (myininaya):

Find the other leg You know Pythagorean thm

myininaya (myininaya):

\[a^2+(-2)^2=3^2 \] Solve for a to find the measurement of the opposite side of angle theta

myininaya (myininaya):

Recall the following: \[\tan(\theta)=\frac{opp}{adj} , \sin(\theta)=\frac{opp}{hyp}\]

OpenStudy (anonymous):

a = √5 yes?

myininaya (myininaya):

\[\sin(\theta)=?\]

myininaya (myininaya):

oops one more thing that y value is below the x-axis so we actually have